OpenStudy (anonymous):

sec theta=1/(?) a. tan theta b. sin theta c. csc theta d. cos theta

OpenStudy (anonymous):

sec theta = 1/cos theta

OpenStudy (anonymous):

\[\frac{1}{\cos \theta} = \sec \theta\] is a fundamental identity in trig, you should remember it